# Uncertainty Quantification for Language Models: A Suite of Black-Box, White-Box, LLM Judge, and Ensemble Scorers

### v3

@@ 3.3 White-Box UQ Scorers
$LNTP(y_{i})=\prod_{t\in y_{i}}p_{t}^{\frac{1}{L_{i}}},$ (10)
@@ 3.4 LLM-as-a-Judge Scorers
We employ LLM-as-a-Judge as an additional method for obtaining response-level confidence scores. In this approach, we concatenate a question-response pair and pass it to an LLM with a carefully constructed instruction prompt that directs the model to evaluate the correctness of the response. We adapt our instruction prompt from Xiong et al. (2024), instructing the LLM to score responses on a 0-100 scale, where a higher score indicates a greater certainty that the provided response is correct. These scores are then normalized to a 0-to-1 scale to maintain consistency with our other confidence scoring methods. The complete prompt template is provided in Appendix A.
@@ 3.5 Ensemble Scorer
$\hat{s}(y_{i};\tilde{\mathbf{y}}_{i},\mathbf{w})=\sum_{k=1}^{K}w_{k}\hat{s}_{k}(y_{i};\tilde{\mathbf{y}}_{i}),$ (12)
@@
Tuning the ensemble requires a sample of LLM responses $y_{1},...,y_{n}$ to a set of $n$ prompts, a set of $K$ confidence scores for each response $\{s_{1}(y_{i};\tilde{\mathbf{y}}_{i}),...,s_{K}(y_{i};\tilde{\mathbf{y}}_{i})\}_{i=1}^{N}$, and corresponding binary hallucination indicators $h(y_{1}),...,h(y_{n})$. [^8] Note that $h(y_{1}),...,h(y_{n})$ serve as ‘ground truth’ labels in the classification objective function. Given a classification objective function, the ensemble weights $\mathbf{w}$ can be tuned with an optimization routine. [^9] We use optuna (Akiba et al., 2019) for optimization with default settings (more details available here). If the objective is threshold-agnostic, the weights and threshold $\tau$ can be tuned sequentially. For a threshold-dependent objective, the weights and threshold can be tuned jointly. See Appendix B for more details on ensemble tuning.
@@ 4.1 Experiment Setup
We conduct a series of experiments to assess the hallucination detection performance of the various scorers. To accomplish this, we leverage a set of publicly available benchmark datasets that contain questions and answers. To ensure that our answer format has sufficient variation, we use two benchmarks with numerical answers (GSM8K (Cobbe et al., 2021) and SVAMP (Patel et al., 2021)), two with multiple-choice answers (CSQA (Talmor et al., 2022) and AI2-ARC (Clark et al., 2018)), and two with open-ended text answers (PopQA (Mallen et al., 2023) and NQ-Open (Lee et al., 2019)).
We sample 1000 questions for each of the six benchmarks. For each question, we generate an original response and $m=15$ candidate responses using four LLMs: GPT-3.5 (OpenAI, ), GPT-4o (OpenAI, ), Gemini-1.0-Pro (Google, ), and Gemini-1.5-Flash (Google, ). [^10] We use a large number of candidate responses ($m=15$) to ensure robust comparisons across black-box scorers. In practice, fewer candidates can be used. For an experimental evaluation of the impact of $m$ on performance, refer to Appendix C. Candidate responses are generated using a temperature of 1.0. For each response, we use the corresponding candidate responses generated by the same LLM to compute the full suite of black-box UQ scores. We also compute LLM-as-a-Judge scores for each response using three different judge models: GPT-3.5, GPT-4o, and Gemini-1.5-Flash. [^11] We were unable to use Gemini-1.0-Pro as a judge model because it was retired during the course of our experiments. Lastly, we compute white-box scores for the GPT-4o, Gemini-1.0-Pro, and Gemini-1.5-Flash responses. [^12] Our instance of GPT-3.5 did not support token probability access. Hence, we were unable to compute white-box UQ scores for the GPT-3.5 responses. We evaluate the hallucination detection performance of all individual scorers as well as our ensemble scorer for each of the benchmarks using various metrics. All scores are computed using this paper’s companion toolkit, uqlm. [^13] Using an n1-standard-16 machine (16 vCPU, 8 core, 60 GB memory) with a single NVIDA T4 GPU attached, our experiments took approximately 0.5-3 hours per LLM-dataset combination to complete.
@@ 4.2 Experiment Results
To assess the performance of the scorers as hallucination classifiers, we evaluate the performance of the confidence scores in a threshold-agnostic fashion. [^14] In our experiments, we label ‘correct’ LLM responses as 1 and ‘incorrect’ responses as 0. Under this setting, we use the AUROC-optimized ensemble weights and compute the ensemble’s AUROC using 5-fold cross-validation. The final reported AUROC is obtained by averaging the AUROC values across the five holdout sets.
Model Metric NQ-Open PopQA GSM8K SVAMP CSQA AI2-ARC
Gem.-1.0-Pro
AUROC 0.84 0.86 0.84 0.88 0.87 0.90
Best Scorer Ensemble Ensemble Ensemble NSN GPT-4o Ensemble
Gem.-1.5-Flash
AUROC 0.79 0.87 0.82 0.89 0.79 0.93
Best Scorer Ensemble Ensemble Ensemble MTP Ensemble LNTP
GPT-3.5
AUROC 0.76 0.72 0.84 0.88 0.84 0.91
Best Scorer Ensemble NCS Ensemble Ensemble Ensemble Ensemble
GPT-4o
AUROC 0.73 0.91 0.90 0.89 0.84 0.86
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble LNTP
Figure 1 presents the AUROC scores for all scorers across the 24 LLM-dataset scenarios, while Table 1 highlights the best-performing scorer for each scenario. The AUROC values for the scenario-specific best scorers range from 0.72 for GPT-3.5 responses on PopQA (NCS) to 0.93 for Gemini-1.5-Flash responses on AI2-ARC (LNTP). Overall, the top-performing scorers for each scenario exhibit strong hallucination detection performance, with AUROC values greater than 0.8 for 19 out of 24 scenarios. However, some scorers perform only slightly better than a random classifier in hallucination detection in certain scenarios, such as the GPT-3.5 LLM judge for all four GSM8K scenarios.
@@
When comparing AUROC values across scorers, we find our ensemble scorer outperforms its individual components in 18 out of 24 scenarios, demonstrating the benefits of use-case-specific optimization. The rankings of individual scorers, however, vary significantly across scenarios, with black-box, LLM-as-a-Judge, and white-box methods achieving top non-ensemble performance in 11, 8, and 5 scenarios respectively. Consistent with previous studies (Kuhn et al., 2023; Manakul et al., 2023; Lin et al., 2024; Farquhar et al., 2024), NLI-based scorers (NSN and NCP) typically lead performance among black-box scorers, achieving the highest AUROC in 18 out of 24 scenarios. Among LLM judges, GPT-4o consistently outperforms GPT-3.5 and Gemini-1.5-Flash, ranking highest in 19 out of 24 scenarios. Finally, the two white-box scorers perform approximately equally, with very similar AUROC scores in the vast majority of scenarios.
@@
Figure 2 displays the F1-scores for each scorer, while Table 2 summarizes the precision, recall, and F1 metrics for the top-performing scorer across all 24 evaluation scenarios. Overall, the results are consistent with the AUROC experiments. Again, the ensemble scorer outperforms its individual components in most scenarios, achieving highest F1-score in 16 out of 24 scenarios. As in the AUROC experiments, the NLI-based scorers (NSN and NCP) are consistently the top-performing black-box scorers (19 out of 24 scenarios), GPT-4o is consistently the top-performing judge (22 out of 24 scenarios), and the two white-box scorers perform approximately equally. Interestingly, the strong performance of GPT-4o as an LLM judge is consistent with its performance in answering questions as the original LLM, where it outperforms GPT-3.5, Gemini-1.5-Flash, and Gemini-1.0-Pro in baseline accuracy across all six benchmarks (see Figure 3).
Model Metric NQ-Open PopQA GSM8K SVAMP CSQA AI2-ARC
Gem.-1.0-Pro
Precision 0.54 0.45 0.58 0.86 0.82 0.95
Recall 0.73 0.69 0.70 0.91 0.98 1.00
F1-Score 0.62 0.54 0.63 0.89 0.89 0.97
Best Scorer EMR NSN Ensemble NSN GPT-4o Ensemble
Gem.-1.5-Flash
Precision 0.62 0.55 0.58 0.89 0.84 0.95
Recall 0.75 0.82 0.88 0.98 0.97 1.00
F1-Score 0.67 0.66 0.70 0.93 0.90 0.97
Best Scorer Ensemble GPT-4o NCP NCP Ensemble Ensemble
GPT-3.5
Precision 0.57 0.32 0.54 0.83 0.85 0.94
Recall 0.85 0.79 0.71 0.94 0.96 1.00
F1-Score 0.68 0.42 0.60 0.89 0.90 0.97
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble Ensemble
GPT-4o
Precision 0.67 0.72 0.84 0.95 0.85 0.99
Recall 0.87 0.79 0.88 0.99 0.97 1.00
F1-Score 0.75 0.75 0.85 0.97 0.91 0.99
Best Scorer Ensemble Ensemble Ensemble Ensemble NCP Ensemble
@@
A notable exception to this trend occurs with black-box scorers on Gemini-1.5-Flash responses to CSQA and AI2-ARC datasets, where no accuracy gains are observed despite filtering. This absence of improvement is consistent with the remarkably low AUROC values for these scorers in the same scenarios shown in Figure 1. Our supplemental analysis in Appendix C reveals that this limitation stems from the lack of response diversity in these two scenarios, with average exact match rates of 0.99 and 1.00, respectively, indicating near-uniformity in candidate responses that provide little signal for black-box scorers to leverage.
@@ 5 Discussion
Choosing the right confidence scorer for an LLM system depends on several factors, including API support, latency requirements, LLM behavior, and the availability of graded datasets. If the API supports access to token probabilities in LLM generations, white-box scorers can be implemented without adding latency or generation costs. However, if the API does not provide access to token probabilities, black-box scorers and LLM-as-a-Judge may be the only feasible options. When choosing among black-box and LLM-as-a-Judge scorers, latency requirements are a key consideration. For low-latency applications, practitioners should avoid higher-latency black-box scorers such as NLI-based scorers (NSN and NCP), opting instead for faster black-box scorers or LLM-as-a-Judge. If latency is not a concern, any of the black-box scorers may be suitable.
Black-box scorers may struggle when LLMs exhibit minimal variation in sampled responses, as shown in our experiments (Appendix C). In such cases, white-box scorers or LLM-as-a-Judge approaches are preferable. For LLM-as-a-Judge implementations, our findings reveal that an LLM’s accuracy on a specific dataset positively relates to its ability to judge responses to questions from that same dataset, providing a practical criterion for judge selection. Finally, if a graded dataset is available, practitioners can tune an ensemble of various confidence scores to improve hallucination detection, as detailed in Appendix B. Our experiments demonstrate that a tuned ensemble can potentially provide more accurate confidence scores than individual scorers. By considering these factors and choosing the right confidence score, practitioners can improve the performance of their LLM system.
@@
We note a few important limitations of this work. First, although our experiments leverage six question-answering benchmark datasets spanning three types of questions, we acknowledge that other types of questions exist. For instance, our experiments do not explore summarization or other long-form tasks, and our findings on scorer-specific performance may not generalize to these types of questions. [^15] For practical reasons, we selected benchmark datasets containing questions that could be graded computationally. Second, while we conduct experiments using four LLMs, performance of the various scorers may differ for other LLMs. Note that for different LLMs, differences in token probability distributions will impact the behavior of white-box scorers, and the degree of variation in responses to the same prompt will affect the performance of black-box scorers. Additionally, LLM-as-a-Judge may have better performance using more recently released, higher-performance LLMs or alternative instruction prompts. Lastly, we note that we have only considered linear ensembles in our experiments. For future work, we suggest deeper explorations of ensembling techniques and their impact on hallucination detection performance.

### v4

@@ 3.3 White-Box UQ Scorers
$LNTP(y_{i};x_{i})=\prod_{t\in y_{i}}p_{t}^{\frac{1}{L_{i}}},$ (10)
@@ 3.4 LLM-as-a-Judge Scorers
We employ an LLM-as-a-Judge scorer $J(y_{i};x_{i})$ as an additional method for obtaining response-level confidence scores. In this approach, we concatenate a question-response pair and pass it to an LLM with a carefully constructed instruction prompt that directs the model to evaluate the correctness of the response. We adapt our instruction prompt from xiong2024llmsexpressuncertaintyempirical, instructing the LLM to score responses on a 0-100 scale, where a higher score indicates a greater certainty that the provided response is correct. These scores are then normalized to a 0-to-1 scale to maintain consistency with our other confidence scoring methods. The complete prompt template is provided in Appendix A.
@@ 3.5 Ensemble Scorer
$\hat{s}(y_{i};\tilde{\mathbf{y}}_{i},x_{i},\mathbf{w})=\sum_{k=1}^{K}w_{k}\hat{s}_{k}(y_{i};\tilde{\mathbf{y}}_{i},x_{i}),$ (12)
@@
Tuning the ensemble requires a sample of LLM responses $y_{1},...,y_{n}$ to a set of $n$ prompts, a set of $K$ confidence scores for each response $\{s_{1}(y_{i};\tilde{\mathbf{y}}_{i},x_{i}),...,s_{K}(y_{i};\tilde{\mathbf{y}}_{i},x_{i})\}_{i=1}^{N}$, and corresponding binary hallucination indicators $h(y_{1};y_{1}^{*},x_{1}),...,h(y_{n};y_{n}^{*},x_{n})$. [^11] Note that $h(y_{1};y_{1}^{*},x_{1}),...,h(y_{n};y_{n}^{*},x_{n})$ serve as ‘ground truth’ labels in the classification objective function. Given a classification objective function, the ensemble weights $\mathbf{w}$ can be tuned with an optimization routine. [^12] We use optuna (Akiba_Optuna_A_next-generation_2019) for optimization with default settings (more details available here). If the objective is threshold-agnostic, the weights and threshold $\tau$ can be tuned sequentially. For a threshold-dependent objective, the weights and threshold can be tuned jointly. Because scorer performance depends on both the dataset and the underlying LLM, the ensemble weights are tuned per use case (chosen LLM and dataset), and the resulting weights are intended for in-domain deployment. See Appendix B for more details on ensemble tuning.
@@ 4.1 Experiment Setup
We conduct a series of experiments to assess the hallucination detection performance of the various scorers. To accomplish this, we leverage a set of publicly available benchmark datasets that contain questions and answers. To ensure that our answer format has sufficient variation, we use two benchmarks with numerical answers (GSM8K (cobbe2021trainingverifierssolvemath) and SVAMP (patel2021nlpmodelsreallyable)), two with multiple-choice answers (CSQA (talmor2022commonsenseqa20exposinglimits) and AI2-ARC (clark2018thinksolvedquestionanswering)), and two with open-ended text answers (PopQA (mallen-etal-2023-trust) and NQ-Open (lee2019latentretrievalweaklysupervised)).
We sample 1000 questions for each of the six benchmarks. For each question, we generate an original response and $m=15$ candidate responses using four LLMs: GPT-4o (OpenAI_doc), GPT-4o-mini (OpenAI_doc), Gemini-2.5-Flash (gemini_doc), and Gemini-2.5-Flash-Lite (gemini_doc). [^13] We use a large number of candidate responses ($m=15$) to ensure robust comparisons across black-box scorers. In practice, fewer candidates can be used. For an experimental evaluation of the impact of $m$ on performance, refer to Appendix D. Candidate responses are generated using a temperature of 1.0. For each response, we use the corresponding candidate responses generated by the same LLM to compute the full suite of black-box UQ scores. We extract the log-probabilities from the generated responses to compute white-box UQ scores. Lastly, we compute LLM-as-a-Judge scores for each response using the same four LLMs. We evaluate the hallucination detection performance of all individual scorers as well as our ensemble scorer for each of the benchmarks using various metrics, with grading processes detailed in Appendix C. All scores are computed using this paper’s companion toolkit, uqlm. [^14] Using an n1-standard-16 machine (16 vCPU, 8 core, 60 GB memory) with a single NVIDA T4 GPU attached, our experiments took approximately 0.5-3 hours per LLM-dataset combination to complete.
@@ 4.2 Experiment Results
To assess the performance of the scorers as hallucination classifiers, we evaluate the performance of the confidence scores in a threshold-agnostic fashion. [^15] In our experiments, we label ‘correct’ LLM responses as 1 and ‘incorrect’ responses as 0. Under this setting, we use the AUROC-optimized ensemble weights and compute the ensemble’s AUROC using 5-fold cross-validation. [^16] Scorer performance depends on both the dataset and the LLM, so we tune the ensemble for a specific LLM–dataset pair. Our experiments reflect this in-domain setup; we do not evaluate out-of-distribution generalization across datasets or domains. The final reported AUROC is obtained by averaging the AUROC values across the five holdout sets.
Figure 1 presents the AUROC scores for all scorers across the 24 LLM-dataset scenarios, while Table 1 highlights the best-performing scorer for each scenario. The AUROC values for the scenario-specific best scorers range from 0.729 for GPT-4o responses on NQ-Open (Ensemble) to 0.986 for Gemini-2.5-Flash-Lite responses on GSM8K (Ensemble). Overall, the top-performing scorers for each scenario exhibit strong hallucination detection performance, with AUROC values greater than 0.8 for 19 out of 24 scenarios. However, some scorers perform no better than or only negligibly better than a random classifier in hallucination detection in certain scenarios, such as the GPT-4o-mini LLM judge for all four GSM8K scenarios.
@@
When comparing AUROC values across scorers, we find our ensemble scorer outperforms its individual components in 20 out of 24 scenarios, demonstrating the benefits of use-case-specific optimization. The rankings of individual scorers, however, vary significantly across scenarios, with LLM-as-a-Judge, black-box, and white-box methods achieving top non-ensemble performance in 11, 7, and 6 scenarios respectively. NLI-based scorers (NSN and NCP) exhibit strong performance, achieving the highest AUROC in 13 out of 24 scenarios among black-box scorers. Among LLM judges, GPT-4o and Gemini-2.5-Flash consistently outperform their smaller counterparts (GPT-4o-mini and Gemini-2.5-Flash-Lite), with GPT-4o ranking as the highest-performing judge in 13 out of 24 scenarios and Gemini-2.5-Flash in 10 out of 24 scenarios. Notably, Gemini-2.5-Flash tended to be the highest performing LLM judge on the math benchmarks (6 out of 8 scenarios), and GPT-4o tended to be the highest performing judge on the short-answer benchmarks (6 out of 8 scenarios). Finally, the two white-box scorers perform approximately equally, with very similar AUROC scores in the vast majority of scenarios.
Model Metric NQ-Open PopQA GSM8K SVAMP CSQA AI2-ARC
Gem.-2.5-Flash AUROC 0.749 0.826 0.771 0.854 0.800 0.975
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble LNTP
Gem.-2.5-Flash-Lite AUROC 0.806 0.880 0.986 0.968 0.840 0.965
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble LNTP
GPT-4o AUROC 0.729 0.798 0.979 0.938 0.844 0.860
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble LNTP
GPT-4o-Mini AUROC 0.778 0.880 0.982 0.930 0.822 0.935
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble LNTP
@@
Model Metric NQ-Open PopQA GSM8K SVAMP CSQA AI2-ARC
Gem.-2.5-Flash Precision 0.597 0.688 0.950 0.968 0.856 0.985
Recall 0.907 0.890 0.999 0.999 0.986 0.987
F1-Score 0.718 0.774 0.974 0.983 0.916 0.986
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble NSN
Gem.-2.5-Fl.-Lt. Precision 0.587 0.680 0.960 0.977 0.863 0.968
Recall 0.838 0.864 0.980 0.976 0.973 0.997
F1-Score 0.688 0.759 0.969 0.977 0.915 0.982
Best Scorer Ensemble Ensemble Ensemble Gem.-2.5-Fl. Ensemble Gem.-2.5-Fl.
GPT-4o Precision 0.649 0.727 0.958 0.978 0.853 0.987
Recall 0.917 0.880 0.987 0.990 0.972 1.000
F1-Score 0.756 0.795 0.972 0.984 0.909 0.993
Best Scorer Ensemble Ensemble Ensemble Gem.-2.5-Fl. NCP Gem.-2.5-Fl.
GPT-4o-Mini Precision 0.616 0.700 0.906 0.969 0.840 0.952
Recall 0.867 0.876 0.978 0.988 0.991 0.998
F1-Score 0.718 0.776 0.940 0.978 0.909 0.975
Best Scorer Ensemble Ensemble Ensemble Ensemble Ensemble Gem.-2.5-Fl.
Figure 2 displays the F1-scores for each scorer, while Table 2 summarizes the precision, recall, and F1 metrics for the top-performing scorer across all 24 evaluation scenarios. The results are largely consistent with the AUROC experiments. The ensemble scorer outperforms its individual components in most scenarios, achieving highest F1-score in 17 out of 24 scenarios. Similar to the AUROC experiments, the NLI-based scorers (NSN and NCP) often outperform other black-box scorers (18 out of 24 scenarios), consistent with the findings of previous studies (kuhn2023semanticuncertaintylinguisticinvariances; manakul2023selfcheckgptzeroresourceblackboxhallucination; lin2024generatingconfidenceuncertaintyquantification; Farquhar2024). Once again, the Gemini-2.5-Flash judge outperforms other LLM judges on the math benchmarks (8 out of 8 scenarios), while the GPT-4o judge outperforms other LLM judges on the short answer benchmarks (6 out of 8 scenarios). Interestingly, the strong LLM judge performance of Gemini-2.5-Flash on the math benchmarks and GPT-4o on the short answer benchmarks are consistent with their respective strengths when generating answers as the original LLM, where they achieved highest accuracy among the four LLMs on those same benchmarks (see Figure 3). Lastly, the two white-box scorers perform approximately equally, as in the AUROC experiments.
@@ 5 Discussion
Choosing the right confidence scorer for an LLM system depends on several factors, including API support, latency requirements, LLM behavior, and the availability of graded datasets. If the API supports access to token probabilities in LLM generations, white-box scorers can be implemented without adding latency or generation costs. [^17] The white-box scorers we consider in this work require only a single generation per prompt. However, sampling-based white-box methods also exist. See, for example, kuhn2023semanticuncertaintylinguisticinvariances; scalena2025eagerentropyawaregenerationadaptive; vashurin2025uncertaintyquantificationllmsminimum; Vashurin_2025; qiu2024semanticdensityuncertaintyquantification. However, if the API does not provide access to token probabilities, black-box scorers and LLM-as-a-Judge may be the only feasible options. When choosing among black-box and LLM-as-a-Judge scorers, latency requirements are a key consideration. For low-latency applications, practitioners should avoid higher-latency black-box scorers such as NLI-based scorers (NSN and NCP), opting instead for faster black-box scorers or LLM-as-a-Judge. If latency is not a concern, any of the black-box scorers may be suitable.
For LLM-as-a-Judge implementations, our findings reveal that an LLM’s accuracy on a specific dataset positively relates to its ability to judge responses to questions from that same dataset, providing a practical criterion for judge selection. Finally, if a graded dataset is available, practitioners can tune an ensemble of various confidence scores to improve hallucination detection, as detailed in Appendix B. Our experiments demonstrate that a tuned ensemble can potentially provide more accurate confidence scores than individual scorers. By considering these factors and choosing the right confidence score, practitioners can improve the performance of their LLM system.
@@
We note two important ethical considerations related to using confidence scores. First, a confidence score reflects model uncertainty rather than ground-truth-based correctness, and high scores can induce over-reliance in high-risk settings. We therefore caution against using confidence scores as decision guarantees in domains such as medicine, law, or finance without external human review and domain-specific safeguards. Second, we encourage monitoring for distributional fairness and content diversity, including per-segment error-rate evaluation, and, where applicable, safeguards that prevent systematic removal of accurate responses for particular groups.
@@
We note a few important limitations of this work. First, although our experiments leverage six question-answering benchmark datasets spanning three types of questions, we acknowledge that our findings on scorer-specific performance may not generalize to other types of questions. [^18] For practical reasons, we selected benchmark datasets containing questions that could be graded computationally. For example, long-form generation (such as summarization) can mix true and false claims within a single output and may require claim-level decomposition to obtain useful uncertainty signals. Likewise, correctness in code generation depends on syntax and execution, so black-box consistency measurements will likely require alternative methods.
Second, while we conduct experiments using four LLMs, performance of the various scorers may differ for other LLMs. Note that for different LLMs, differences in token probability distributions will impact the behavior of white-box scorers, and the degree of variation in responses to the same prompt will affect the performance of black-box scorers. LLM-as-a-Judge performance may vary significantly depending on the choice of LLM and instruction prompts used.
Additional limitations relate to the ensemble approach. First, while the ensemble approach is designed to optimize in-domain performance, we have not evaluated out-of-distribution generalization, including cross-dataset transfer of learned weights; we leave this to future work. Second, our experiments consider only linear ensembles. Future work should examine non-linear ensembling and its impact on hallucination detection, e.g., monotonic generalized additive models that retain interpretability while capturing modest non-linear effects, mixture-of-experts, and tree-based ensembles. These may outperform linear averaging when scorer signals interact or exhibit non-linear relationships. Practical trade-offs to study include latency, overfitting risk on small graded sets, and the need for interpretability in high-stakes deployments. Third, tuning the ensemble requires a graded dataset. For tasks with trivial grading (e.g., arithmetic, multiple-choice, short-answer), where there are well-defined, easy-to-recognize correct answers and comparisons with generated text can be computed automatically, this can be done by sampling a representative set of questions, generating responses, computing the selected UQ scores, and using the automatically derived labels to fit the weights. For tasks with non-trivial grading (e.g., summarization), practitioners can begin with a small human-graded dataset (e.g., a few hundred items); additional labels can then be added incrementally from production logs to further optimize the weights and improve ensemble performance.